OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <8, 24> Neither Parallel Orthogonal

OpenStudy (anonymous):

take dot product u dot v, if orthogonal it will = 0, if there is some scalar lambda that you can multiply by u to get v then its parallel, if neither, then neither.

OpenStudy (anonymous):

scalar lamda?

OpenStudy (anonymous):

you can say scalar S, i dont know how to type lambda in here lol

OpenStudy (anonymous):

is there a number that you can multiply by both components of u to get the components of v?

OpenStudy (anonymous):

so -2 times what equals 24?

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

correct, and does -12 * 6 = 8?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then its not parallel

OpenStudy (anonymous):

so do the dot product of u and v and if its 0 then its orthogonal if not then its neither, since we know its not parallel

OpenStudy (anonymous):

@cstang, do you know how to do u dot v?

OpenStudy (anonymous):

i got 0 for the dot product

OpenStudy (anonymous):

6·8 + (-2) ·24 = 48+(-48) = 0